# Coupled gas and bubble dynamics at the solidification front

### v1

@@ I Introduction
$\frac{C_{L}(x,t)}{C_{0}}=1+\left(\frac{1-k}{k}\right)\left[\exp\left(-\frac{V_{sf}}{D}x\right)-\exp\left(-\frac{V_{sl}(1-k)\left(x+kV_{sf}t\right)}{D}\right)\right]$ (1)
where $t$ is the time, $D$ is the diffusion coefficient of the gas, $C_{L}$ is the gas concentration at a distance $x$ from the solidification front and $C_{0}$ is the bulk gas concentration in the liquid far from the solidification front. The partition coefficient $k$ is related to the solubility of the gas in the ice and is defined as the ratio of gas solubility in the solid phase to the one in the liquid phase. The local segregation of solutes is especially critical for examining the morphology of the solidification front, the occurrence of constitutional supercooling leading to the formation of premelted liquid films [18, 19], as well as for understanding the behavior of objects in proximity to the front [13].
@@ II.1 Materials and sample preparation
An aqueous fluorophore, Sulforhodamine B, was obtained from FluoTechnik, alongside a Sodastream Terra device to facilitate the injection of carbon dioxide into water. The solutions were prepared by dissolving Sulforhodamine B in distilled water, followed by stirring for approximately one minute, to obtain a $10^{-5}M$ solution. Carbon dioxide was introduced into the solution by administering three one-second pulses using the Sodastream Terra device. Subsequently, the solution was immediately introduced into a rectangular Hele-Shaw cell via capillary action. This cell comprised two glass slides (Menzel, thickness 0,130,19mm0,13-0,19mm), held apart by two spacers composed of double-sided adhesive tape, which maintained a consistent sample thickness of approximately $100\mu m$. The cell was sealed on one side with a plastic support (used to pull the sample through the temperature gradient) and on the opposite side with nail polish to prevent both evaporation and leakage.
@@ II.2 Experimental setup and typical solidification parameters
Unidirectional solidification experiments were conducted utilizing the experimental setup described previously [19] and visualized through confocal fluorescence microscopy (Figure 1). The Hele-Shaw cell was positioned beneath the microscope objective on two Peltier modules and was pulled through the temperature gradient at a constant velocity by a stepper motor (Micos Pollux Drive stepper motor with VT-80 translation stage (PI, USA)). The temperature gradient was established by two Peltier modules and managed with a TEC-1122 Dual Thermo Electric Cooling Temperature Controller from Meerstetter Engineering, Switzerland. The stage was enhanced by enclosing it within a Plexiglas box supplied with dry air to minimize humidity (typically between 5 and 10%) to avert condensation and freezing on cold surfaces. This measure prevents potential friction with the sample and mitigates light scattering caused by frost.
@@ III.1 Experimental observations
We conducted horizontal solidification experiments by displacing a Hele-Shaw cell filled with carbonated water, at varying freezing velocities $V_{sf}$. A representative series of time-lapse 2D confocal images illustrating the freezing process is presented in figure 2A for $V=20\mu m/s$.
@@
In the observational frame, the solidification front remains stationary, whereas in the sample’s reference frame, the solidification front progresses through the medium at a velocity $V_{sf}$. The ice appears black and the aqueous solution pink. Bubbles nucleate near or at the solidification front within several seconds of freezing. This process is attributed to the segregation of gas at the front, which results in a localized increase in gas concentration surpassing the critical nucleation threshold under the given experimental conditions. The following images, from $t=9s$ to $t=12s$, demonstrate that bubbles can experience an increase in size as a consequence of gas segregation and diffusion or through the coalescence of adjacent bubbles. The transfer of gas between a bubble and the surrounding liquid is regulated by the Laplace pressure, which is a function of the bubble’s curvature. The bubbles eventually interact with the solidification front and undergo deformation during their engulfment ($t=16s$ and $t=33s$). Mechanical stresses acting on bubbles arise from the pushing effect of the solidification front. This mechanism has been described for solid particles [44], and bubbles experience a similar effect [45]. After engulfment of the bubbles by the front ($t=44s$), another cycle of gas segregation is initiated ($t=57s$), until the next nucleation event.
@@
Figure 2: Representative images obtained through confocal cryomicroscopy. These time-lapse sequences of typical two-dimensional confocal images illustrate the freezing process of carbonated water. In each image, the lower, darker region corresponds to the ice phase, whereas the upper, lighter region corresponds to the liquid phase. Bubbles, being non-fluorescent, appear as dark regions within the liquid. (A) The time-lapse sequences display typical 2D cryo-confocal images demonstrating the nucleation, interaction, and engulfment of bubbles during the solidification process. Velocity of the solidification front $V_{sf}=20\mu m/s$. The bubble formation in these images is ascribed to nucleation induced by gas accumulation at the advancing solidification front, attributed to gas segregation. (B), (C) The sequences also illustrate the formation and evolution of cylindrical bubbles at the solidification front during solidification for $V_{sf}=20\mu m/s$ (B) and $V_{sf}=1\mu m/s$ (C).
@@
The deformation of these cylindrical bubbles and the influence of solidification velocity, gas concentration, pressure, and mass transport of gas have been thoroughly examined in previous studies [7, 32, 38]. Moreover, models have been formulated to predict the formation of such bubbles [41, 40]. Although a sharp transition from circular to cylindrical bubbles was predicted [40], we did not observe it here. In our steady state regime, with a constant front velocity and constant temperature gradient, there is a continuous transition from circular to cylindrical bubbles. The stability, and thus the time that cylindrical bubbles remains at the front, increases as the front velocity decreases. Cylindrical bubbles persist longer (up to $1000s$ and more) at reduced solidification velocities ($1-2\mu m/s$, with some observed to grow laterally when the solidification velocity $V_{sf}$ is below $5\mu m/s$ (Figure 2C). Conversely, at elevated solidification velocities $V_{sf}$ ($5-20\mu m/s$), cylindrical bubbles may still form; however, they are smaller and become engulfed by the solidification front shortly after formation (Figure 2B). This is attributed to the inadequate gas diffusion from the surrounding liquid to the exposed tip of the cylindrical bubbles at high solidification velocities.
@@ III.2 Average nucleation rates
We first examine the impact of solidification velocity $V_{sf}$ on the average rate of bubble nucleation, quantitatively expressed as the number of bubbles formed per second (Figure 3). This analysis currently focuses for now solely on the incidence of bubble nucleation events, deliberately excluding any subsequent phenomena such as changes in bubble size, growth kinetics, or interactions with adjacent bubbles. The freezing front velocity $V_{sf}$ is systematically increased from $1\mu m/s$ to $20\mu m/s$. At each specified velocity, the solidification process continues long enough to collect a sufficient number of bubbles, typically between 20 and 70 for velocities in the range of $8-20\mu m/s$ . At lower velocities, the number of bubbles remained very limited—typically fewer than 10 because of the small number of bubble nucleation. By using a single sample with a gradually increasing or decreasing solidification velocity, it is ensured that any observed variations in nucleation behavior are attributed to velocity changes rather than fluctuations in the initial gas concentration. Additionally, an alternative set of measurements is conducted by reducing the velocity from $20\mu m/s$ to $1\mu m/s$.
We observe that the average nucleation rate increases with the solidification velocity for both the increasing or decreasing ramps. The enhanced nucleation rate observed at higher solidification velocities is consistent with an increased segregation of dissolved gases at the solidification front. Faster solidification front velocities accelerate local gas enrichment, making it more likely for the critical nucleation concentration $C^{*}$ to be reached within a shorter timescale.
Furthermore, nucleation rates demonstrate consistency and reproducibility across both velocity ramp protocols. A notable discrepancy is observed for $V=20\mu m/s$, wherein the nucleation rate observed during the decreasing velocity ramp is markedly lower than that in the increasing velocity ramp. This inconsistency is likely due to the transient period necessary for the system to attain a steady state, coupled with the segregation of dissolved gas at the solidification front and the attainment of the critical nucleation concentration $C^{*}$. In the experiments involving increasing velocities, the gas segregation is pre-established from the onset when the velocity transitions from $18\mu m/s$ to $20\mu m/s$. Conversely, in the decreasing velocity ramps, it is probable that the steady state is not achieved for the initial velocity examined ($20\mu m/s$).
@@
Figure 3: Evolution of the nucleation rate of bubbles as a function of the velocity of solidification $V_{sf}$. Measurement were done with similar sample and identical experimental conditions but with either an increasing or a decreasing solidification front velocity.
@@ III.3 Dynamics of bubble nucleation
We now focus on the dynamics of bubble nucleation as influenced by $V_{sf}$ under experimental conditions where the solidification velocity $V_{sf}$ is maintained constant throughout the entirety of a given sample. In each captured image, the count of bubble nucleation events occurring within the field of view is recorded, and their cumulative total is evaluated over the duration of solidification for various solidification velocities (Figure 4).
The slopes of these curves represent the average nucleation rate of gas bubbles for each specified solidification velocity $V_{sf}$, which concurs with the data presented in figure 3. Repeated experiments performed at solidification velocities of $10\mu m/s$ and $15\mu m/s$ consistently yielded convergent results, thereby verifying the reproducibility of the observed nucleation dynamics under controlled conditions. These findings confirm the reliability of the protocol employed for sample preparation and freezing. The rate of nucleation increases with the solidification velocity $V_{sf}$. For example, at $V_{sf}=5\mu m/s$, 30 bubbles nucleated after 2968s, in contrast to 166 bubbles after 2430s for $V_{sf}=10\mu m/s$, 186 bubbles after 930s for $V_{sf}=15\mu m/s$, and 161 bubbles after 600s for $V_{sf}=20\mu m/s$.
@@
Figure 4: Cumulative number of nucleated bubbles as a function of solidification duration under varying velocities. Multiple experiments were performed at solidification velocities of $V_{sf}=10\mu m/s$ and $V_{sf}=15\mu m/s$ to validate the reproducibility of the measurements. The plateau length indicates the lag period $t_{plateau}$ between successive bubble nucleation events.
We now re-examine the data presented in figure 4 to determine the total proportion of the field of view occupied by the bubbles in the liquid region. This analysis, depicted in figure 5, allows for the differentiation between the step of bubble nucleation and growth—resulting in an increase in the area occupied by the bubbles—and the subsequent step where they are engulfed by the front, leading to a reduction in the interfacial area. This method utilizes 2D fluorescence microscopy images, applying a threshold to identify and quantify the area occupied by the bubbles in the liquid region. The results, as a function of the duration of directional solidification corresponding to solidification velocities $V_{sf}=5\mu m/s$ and $10\mu m/s$, are presented in figure 5.
@@
At all velocities, successive peaks of varying amplitudes are observed, characterized by rapid expansions of the space occupied by bubbles, followed by a slower decrease. Notably, the onset of the growth phase for each peak aligns with the increasing segments of the time evolution of bubble numbers as depicted in figure 4. These ascending sections of the peaks are attributed to the quasi-simultaneous nucleation of bubbles and their subsequent growth via gas diffusion from the supersaturated liquid. Conversely, the descending segments correspond to the gradual engulfment of these bubbles by the advancing solidification front, resulting in a reduction of their apparent area within the liquid. These curves are significant as they reflect the total area occupied by bubbles in the liquid. Thus, even when bubble growth is dominant, some bubbles may already start being engulfed by the solidification front. Conversely, when the area decreases, indicating that most bubbles are being engulfed, some bubbles may continue to grow simultaneously, even during their engulfment (as is the case for cylindrical bubbles, for example). The peak’s maximum represents the transition point where the bubble growth rate is surpassed by the rate of their engulfment by the solidification front. The plateaus depicted in figure 5 denote not only a period of gas segregation at the solidification front but also a phase during which the majority of the space occupied by the bubbles is engulfed by the advancing solidification front. The characteristic lag time, $t_{plateau}$ (Figure 4), can thus be interpreted as the cumulative duration of bubble nucleation and growth $t_{n}$ (ascending part of the peaks), the bubble engulfment period $t_{e}$ (descending part of the peaks), and a subsequent interval of gas segregation without significant nucleation or engulfment activity.
@@
Furthermore, the minor oscillations observed between the main peaks (Figure 5B) are primarily ascribed to the existence of cylindrical bubbles at the solidification front. In these instances, a portion of the bubble remains emerged in the liquid, sustained by the supply of dissolved gas. These cylindrical bubbles may achieve a quasi-equilibrium state with a consistent width, although the diameter of the emerged tip can fluctuate based on pressure, gas diffusion, and the velocity of solidification. Consequently, in the latter scenario, the area occupied by the emerged tip of the bubble evolves over time. As illustrated in figure 2B and figure 2C, cylindrical bubbles persist longer at the interface and display larger diameters at reduced solidification velocities. Under these circumstances, their extended presence at the interface leads to a greater consumption of gas, as the diffusion rate of gas sufficiently compensates for the engulfment by the advancing solidification front. This gas diffusion towards cylindrical bubbles may consequently decelerate the rise in gas concentration at the solidification front. Thus, we assume that the presence of cylindrical bubbles influences bubble nucleation dynamics, particularly affecting the bubble nucleation rate during solidification. Nonetheless, the nucleation of a solitary bubble can also occur, contributing to these minor oscillations between the peaks.
@@
Based on figure 5, we extracted the peaks describing the evolution of the area occupied by the bubbles and normalized them (figure 6). This methodology enables an accurate comparison of bubble dynamic profiles and quantifies how variations in solidification velocity influence the overall bubble behavior throughout the solidification process. Our findings reveal that for all investigated solidification rates, the observed peaks and corresponding durations of nucleation, growth, and engulfment periods of bubbles remain consistent for a given solidification velocity. For solidification velocities of $V_{sf}=15\mu m/s$ and $V_{sf}=20\mu m/s$, a notable symmetry emerges, indicating that the duration of bubble nucleation and growth $t_{n}$ matches that of bubble engulfment $t_{e}$. At lower solidification velocities, particularly for $V_{sf}=5\mu m/s$ and $V_{sf}=10\mu m/s$, the peaks exhibit asymmetry due to an extended duration of the bubble engulfment phase as the solidification velocity decreases. The analysis of these peaks facilitates the determination of the timescales linked with the coupled gas and bubble dynamics, which are influenced by the solidification front velocity (figure 7).

### v2

@@ I Introduction
$\frac{C_{L}(x,t)}{C_{0}}=1+\left(\frac{1-k}{k}\right)\left[\exp\left(-\frac{V_{sf}}{D}x\right)-\exp\left(-\frac{V_{sf}(1-k)\left(x+kV_{sf}t\right)}{D}\right)\right]$ (1)
where $V_{sf}$ is the solidification front velocity ($m/s$), $t$ is the time ($s$), $D$ is the diffusion coefficient of the gas ($m^{2}/s$), $C_{L}$ is the gas concentration at a distance $x$ from the solidification front and $C_{0}$ is the bulk gas concentration in the liquid far from the solidification front ($g/L$). The partition coefficient $k$ is related to the solubility of the gas in the ice and is defined as the ratio of gas solubility in the solid phase to the one in the liquid phase (and is therefore dimensionless). The local segregation of solutes is especially critical for examining the morphology of the solidification front, the occurrence of constitutional supercooling leading to the formation of premelted liquid films [18, 19], as well as for understanding the behavior of objects in proximity to the front [13].
@@ II.1 Materials and sample preparation
An aqueous fluorophore, Sulforhodamine B, was obtained from FluoTechnik, alongside a Sodastream Terra device to facilitate the injection of carbon dioxide into water. The solutions were prepared by dissolving Sulforhodamine B in distilled water, followed by stirring for approximately one minute, to obtain a $10^{-5}M$ solution. Carbon dioxide was introduced into the solution by administering three one-second pulses (controlled with a timer) using the Sodastream Terra device. The pH was measured using a calibrated electronic pH meter, and the carbonate hardness (KH) was determined using commercially available KH test strips. Both measurements were used to determined the initial $CO_{2}$ concentration before freezing. The temperature (room temperature) was kept constant for all experiments. Subsequently, the solution was immediately introduced into a rectangular Hele-Shaw cell via capillary action. This cell comprised two glass slides (Menzel, thickness 0,130,19mm0,13-0,19mm), held apart by two spacers composed of double-sided adhesive tape, which maintained a consistent sample thickness of approximately $100\mu m$. The cell was sealed on one side with a plastic support (used to pull the sample through the temperature gradient) and on the opposite side with nail polish to prevent both evaporation and leakage. The reproducibility of the approach was assessed by repeating the sample preparation and running the same experiment several times.
The use of thin Hele-Shaw cell is standard in solidification studies, their benefits and relevance have been reported multiple times, for example in the context of solidification studies of transparent alloys.
@@ II.2 Experimental setup and typical solidification parameters
Unidirectional solidification experiments were conducted utilizing the experimental setup described previously [19] and visualized through confocal fluorescence microscopy (Fig. 1). The Hele-Shaw cell was positioned beneath the microscope objective on two Peltier modules and was pulled through the temperature gradient at a constant velocity by a stepper motor (Micos Pollux Drive stepper motor with VT-80 translation stage (PI, USA)). The temperature gradient was established by two Peltier modules. The temperature is measured through thermocouples placed at the surface of the Peltier and continuously monitored. The temperature gradient is managed with a TEC-1122 Dual Thermo Electric Cooling Temperature Controller from Meerstetter Engineering, Switzerland. This setup enables to control independently the temperature gradient $\Delta T=T_{h}-T_{c}$ created between the Peltier modules and the growth velocity $V_{sf}$ of the freezing front. This ability is critical since, as demonstrated in this paper and in previous studies, many of the phenomena associated to solidification are very sensitive to both parameters. We performed all experiments with $T_{h}=15^{\circ}C$ and $T_{c}=-15^{\circ}C$.
The stage was enhanced by enclosing it within a Plexiglas box supplied with dry air to minimize humidity (typically between 5 and 10%) to avert condensation and freezing on cold surfaces. This measure prevents potential friction with the sample and mitigates light scattering caused by frost.
@@
In a typical experiment, the temperature gradient is kept constant, and the translation stage imposing the growth velocity applies a constant growth rate. The solidification front therefore reaches a steady state as soon as its position within the observation frame is constant. This typically occurs within seconds. All the data reported in the paper was obtained in a steady state regime, where the front velocity $V_{sf}$ is constant.
@@ III.1 Experimental observations
We conducted horizontal solidification experiments by displacing a Hele-Shaw cell filled with carbonated water, at varying freezing velocities $V_{sf}$. A representative series of time-lapse 2D confocal images illustrating the freezing process is presented in Fig. 2A for $V=20\mu m/s$.
@@
In the observational frame, the solidification front remains stationary, whereas in the sample’s reference frame, the solidification front progresses through the medium at a velocity $V_{sf}$. The ice appears black and the aqueous solution pink. Bubbles nucleate near or at the solidification front within several seconds of freezing. This process is attributed to the segregation of gas at the front, which results in a localized increase in gas concentration surpassing the critical nucleation threshold under the given experimental conditions.
We performed preliminary observations by imaging the sample at various depths during solidification, and in particular close to the glass slides, to assess the influence of the glass slide on the bubble nucleation behavior. The glass slide do not appear to be preferential nucleation sites for the bubbles, and bubbles do not appear to stick to the glass slide. However, it is still possible that bubble nucleation occurs onto the glass slides and that bubble nearly instantaneously detach from the glass during the initial burst growth stage. All subsequent experiments were therefore performed in the middle of the sample.
The following images, from $t=9s$ to $t=12s$, demonstrate that bubbles can experience an increase in size as a consequence of gas segregation and diffusion or through the coalescence of adjacent bubbles.
The transfer of gas between a bubble and the surrounding liquid is regulated by the Laplace pressure, which is a function of the bubble’s curvature. The bubbles eventually interact with the solidification front and undergo deformation during their engulfment ($t=16s$ and $t=33s$). Mechanical stresses acting on bubbles arise from the pushing effect of the solidification front. This mechanism has been described for solid particles [44], and bubbles experience a similar effect [45]. After engulfment of the bubbles by the front ($t=44s$), another cycle of gas segregation is initiated ($t=57s$), until the next nucleation event.
@@
Figure 2: Representative images obtained through confocal cryomicroscopy. These time-lapse sequences of typical two-dimensional confocal images illustrate the freezing process of carbonated water and the different stages encountered during a typical solidification experiment. In each image, the lower, darker region corresponds to the ice phase, whereas the upper, lighter region corresponds to the liquid phase. Bubbles, being non-fluorescent, appear as dark regions within the liquid. (A) The time-lapse sequences display typical 2D cryo-confocal images demonstrating the nucleation, interaction, and engulfment of bubbles during the solidification process. Velocity of the solidification front $V_{sf}=20\mu m/s$. The bubble formation in these images is ascribed to nucleation induced by gas accumulation at the advancing solidification front, attributed to gas segregation. (B) The sequences also illustrate the formation and evolution of cylindrical bubbles at the solidification front during solidification for $V_{sf}=1\mu m/s$. For each panel, $t_{0}$ corresponds to the time of the first frame of the sequence shown in the panel.
@@
The deformation of these cylindrical bubbles and the influence of solidification velocity, gas concentration, pressure, and mass transport of gas have been thoroughly examined in previous studies [7, 32, 38]. Moreover, models have been formulated to predict the formation of such bubbles [41, 40]. Although a sharp transition from circular to cylindrical bubbles was predicted [40], we did not observe it here. In our steady state regime, with a constant front velocity and constant temperature gradient, there is a continuous transition from circular to cylindrical bubbles. The stability, and thus the time that cylindrical bubbles remains at the front, increases as the front velocity decreases. Cylindrical bubbles persist longer (up to $1000s$ and more) at reduced solidification velocities ($1-2\mu m/s$, with some observed to grow laterally when the solidification velocity $V_{sf}$ is below $5\mu m/s$ (Fig. 2C). Conversely, at elevated solidification velocities $V_{sf}$ ($5-20\mu m/s$), cylindrical bubbles may still form; however, they are smaller and become engulfed by the solidification front shortly after formation (Fig. 2B). This is attributed to the inadequate gas diffusion from the surrounding liquid to the exposed tip of the cylindrical bubbles at high solidification velocities.
Since the nucleation of bubbles occurs close to the solidification front, this also means there is a limited amount of time (typically a few seconds) during which the bubbles could interact with the glass slides, before being engulfed by the solidification front. Although $CO_{2}$ bubbles are known to wet glass, it seems likely that the rapid kinetics observed here prevent such wetting from occurring. We therefore believe that the glass slides have little influence on the observations reported in this study.
@@ III.2 Average nucleation rates
We first examine the impact of solidification velocity $V_{sf}$ on the average rate of bubble nucleation, quantitatively expressed as the number of bubbles formed per second (Fig. 3). This analysis currently focuses for now solely on the incidence of bubble nucleation events, deliberately excluding any subsequent phenomena such as changes in bubble size, growth kinetics, or interactions with adjacent bubbles. The freezing front velocity $V_{sf}$ is systematically increased from $1\mu m/s$ to $20\mu m/s$ (ramp up). Additionally, an alternative set of measurements is conducted by reducing the velocity from $20\mu m/s$ to $1\mu m/s$ (ramp down).
The ramp up and ramp down procedures are the same. The velocity is held constant at a given value to complete the imaging work and then adjusted to a lower (ramp down) or higher (ramp up) value. When the velocity is changed to a new value, it takes only a few seconds for the front to stabilize around its new steady state position, at which point we start imaging. The solidification process continues long enough to collect sufficient statistics, typically between 20 and 70 for velocities in the range of $8-20\mu m/s$ . At lower velocities, the number of bubbles remained very limited—typically fewer than 10 because of the small number of bubble nucleation. We keep track of the number of bubbles that appear during solidification and thus do not need to define a critical bubble size for nucleation and growth. The liquid being saturated with gas, we systematically observe that the bubble first go through an explosive growth stage. Considering the fast imaging rate used (up to 7.5 frames/s), we are thus able to determine precisely the moment when bubbles appear in the liquid. By using a single sample with a gradually increasing or decreasing solidification velocity, it is ensured that any observed variations in nucleation behavior are attributed to velocity changes rather than fluctuations in the initial gas concentration.
We observe that the average nucleation rate increases with the solidification velocity for both the increasing or decreasing ramps. The enhanced nucleation rate observed at higher solidification velocities is consistent with an increased segregation of dissolved gases at the solidification front. Faster solidification front velocities accelerate local gas enrichment, making it more likely for the critical nucleation concentration $C_{n}^{*}$ to be reached within a shorter timescale.
Furthermore, nucleation rates demonstrate consistency and reproducibility across both velocity ramp protocols. A notable discrepancy is observed for $V=20\mu m/s$, wherein the nucleation rate observed during the decreasing velocity ramp is markedly lower than that in the increasing velocity ramp. This inconsistency is likely due to the transient period necessary for the system to attain a steady state, coupled with the segregation of dissolved gas at the solidification front and the attainment of the critical nucleation concentration $C_{n}^{*}$. In the experiments involving increasing velocities, the gas segregation is pre-established from the onset when the velocity transitions from $18\mu m/s$ to $20\mu m/s$. Conversely, in the decreasing velocity ramps, it is probable that the steady state is not achieved for the initial velocity examined ($20\mu m/s$).
@@
Figure 3: Evolution of the nucleation rate of bubbles as a function of the velocity of solidification $V_{sf}$. Measurement were done with similar sample and identical experimental conditions but with either an increasing or a decreasing solidification front velocity. For each data point in this figure, the motor velocity is set to a specific prescribed value. At this constant motor velocity, the number of observed bubble nucleation events is recorded in order to determine the mean nucleation rate corresponding to that condition. Following each measurement, the motor velocity is adjusted to a new value, and the procedure is repeated to obtain data at different velocity for a unique sample. For each velocity of solidification considered, the measurement time ranged from 3 minutes at the highest velocities of solidification to 9 minutes at the lowest velocities of solidification.
@@ III.3 Dynamics of bubble nucleation
We now focus on the dynamics of bubble nucleation as influenced by $V_{sf}$ under experimental conditions where the solidification velocity $V_{sf}$ is maintained constant throughout the entirety of a given sample. In each captured image, the count of bubble nucleation events occurring within the field of view is recorded, and their cumulative total is evaluated over the duration of solidification for various solidification velocities (Fig. 4).
The slopes of these curves represent the average nucleation rate of gas bubbles for each specified solidification velocity $V_{sf}$, which concurs with the data presented in Fig. 3. Repeated experiments performed at solidification velocities of $10\mu m/s$ and $15\mu m/s$ consistently yielded convergent results, thereby verifying the reproducibility of the observed nucleation dynamics under controlled conditions. These findings confirm the reliability of the protocol employed for sample preparation and freezing. The rate of nucleation increases with the solidification velocity $V_{sf}$. For example, at $V_{sf}=5\mu m/s$, 30 bubbles nucleated after 2968s, in contrast to 166 bubbles after 2430s for $V_{sf}=10\mu m/s$, 186 bubbles after 930s for $V_{sf}=15\mu m/s$, and 161 bubbles after 600s for $V_{sf}=20\mu m/s$.
@@
It should be noted that the typical duration of plateau $t_{plateau}$ (tens to hundreds of seconds) is much longer than the time interval between successive frames (typically 0.13 s). We also systematically observe that as soon as bubbles appear, they go through a very fast growth stage and reach within a few frames. The stepwise characteristic of the dynamics observed in Fig. 4 is therefore not due to spatial or temporal resolution issue (the appearance of the very small bubbles that would not be recorded at sufficiently shorter time scale). Hence, the steps reflect the real-time results of the bubble nucleation.
Figure 4: Cumulative number of nucleated bubbles as a function of solidification duration under varying velocities. Multiple experiments on different samples prepared following the same procedure were performed at solidification velocities of $V_{sf}=10\mu m/s$ and $V_{sf}=15\mu m/s$ to validate the reproducibility of the measurements. The plateau length indicates the lag period $t_{plateau}$ between successive bubble nucleation events.
We now re-examine the data presented in Fig. 4 to determine the total proportion of the field of view occupied by the bubbles in the liquid region. This analysis, depicted in Fig. 5, allows for the differentiation between the step of bubble nucleation and growth—resulting in an increase in the area occupied by the bubbles—and the subsequent step where they are engulfed by the front, leading to a reduction in the interfacial area. This method utilizes 2D fluorescence microscopy images, applying a threshold to identify and quantify the area occupied by the bubbles in the liquid region. The results, as a function of the duration of directional solidification corresponding to solidification velocities $V_{sf}=5\mu m/s$ and $10\mu m/s$, are presented in Fig. 5.
@@
At all velocities, successive peaks of varying amplitudes are observed, characterized by rapid expansions of the space occupied by bubbles, followed by a slower decrease. Notably, the onset of the growth phase for each peak aligns with the increasing segments of the time evolution of bubble numbers as depicted in Fig. 4. These ascending sections of the peaks are attributed to the quasi-simultaneous nucleation of bubbles and their subsequent growth via gas diffusion from the supersaturated liquid. Conversely, the descending segments correspond to the gradual engulfment of these bubbles by the advancing solidification front, resulting in a reduction of their apparent area within the liquid. These curves are significant as they reflect the total area occupied by bubbles in the liquid. Thus, even when bubble growth is dominant, some bubbles may already start being engulfed by the solidification front. Conversely, when the area decreases, indicating that most bubbles are being engulfed, some bubbles may continue to grow simultaneously, even during their engulfment (as is the case for cylindrical bubbles, for example). The peak’s maximum represents the transition point where the bubble growth rate is surpassed by the rate of their engulfment by the solidification front. The plateaus depicted in Fig. 5 denote not only a period of gas segregation at the solidification front but also a phase during which the majority of the space occupied by the bubbles is engulfed by the advancing solidification front. The characteristic lag time, $t_{plateau}$ (Fig. 4), can thus be interpreted as the cumulative duration of bubble nucleation and growth $t_{n}$ (ascending part of the peaks), the bubble engulfment period $t_{e}$ (descending part of the peaks), and a subsequent interval of gas segregation without significant nucleation or engulfment activity.
@@
Furthermore, the minor oscillations observed between the main peaks (Fig. 5B) are primarily ascribed to the existence of cylindrical bubbles at the solidification front. In these instances, a portion of the bubble remains emerged in the liquid, sustained by the supply of dissolved gas. These cylindrical bubbles may achieve a quasi-equilibrium state with a consistent width, although the diameter of the emerged tip can fluctuate based on pressure, gas diffusion, and the velocity of solidification. Consequently, in the latter scenario, the area occupied by the emerged tip of the bubble evolves over time. As illustrated in Fig. 2B and Fig. 2C, cylindrical bubbles persist longer at the interface and display larger diameters at reduced solidification velocities. Under these circumstances, their extended presence at the interface leads to a greater consumption of gas, as the diffusion rate of gas sufficiently compensates for the engulfment by the advancing solidification front. This gas diffusion towards cylindrical bubbles may consequently decelerate the rise in gas concentration at the solidification front. Thus, we assume that the presence of cylindrical bubbles influences bubble nucleation dynamics, particularly affecting the bubble nucleation rate during solidification. Nonetheless, the nucleation of a solitary bubble can also occur, contributing to these minor oscillations between the peaks.
@@
Based on Fig. 5, we extracted the peaks describing the evolution of the area occupied by the bubbles and normalized them (Fig. 6). This methodology enables an accurate comparison of bubble dynamic profiles and quantifies how variations in solidification velocity influence the overall bubble behavior throughout the solidification process. Our findings reveal that for all investigated solidification rates, the observed peaks and corresponding durations of nucleation, growth, and engulfment periods of bubbles remain consistent for a given solidification velocity. For solidification velocities of $V_{sf}=15\mu m/s$ and $V_{sf}=20\mu m/s$, a notable symmetry emerges, indicating that the duration of bubble nucleation and growth $t_{n}$ matches that of bubble engulfment $t_{e}$. At lower solidification velocities, particularly for $V_{sf}=5\mu m/s$ and $V_{sf}=10\mu m/s$, the peaks exhibit asymmetry due to an extended duration of the bubble engulfment phase as the solidification velocity decreases. The analysis of these peaks facilitates the determination of the timescales linked with the coupled gas and bubble dynamics, which are influenced by the solidification front velocity (Fig. 7).
@@
The impact of the coalescence of adjacent bubbles should also be briefly discussed. It is well-known that there is a critical size (for a given front velocity) below which bubbles are repelled by the front, and above with they engulfed by the front [45]. One could therefore imagine a situation where small bubbles (below the critical size) would coalesce and therefore be engulfed by the front. In such situation, bubbles coalescence could impact the results reported here. However, we systematically observed rapid growth of bubbles and quasi-systematic engulfment of the bubbles by the front. In other words, bubbles here are almost never small enough to be repelled by the front. We are therefore never in a scenario where the coalescence of bubbles would impact their interaction with the solidification front.